OpenStudy (anonymous):

find the end behavior of f(x)=e^(x-3)-3

OpenStudy (anonymous):

I think you have to take the natural log :)

OpenStudy (solomonzelman):

no

OpenStudy (solomonzelman):

At first, you have a shift ( from \(\large\color{slate}{ f(x)=e^{x} }\) ) 3 units right \(\large\color{slate}{ f(x)=e^{x\color{blue}{-3}} }\) 3 units down, \(\large\color{slate}{ f(x)=e^{x-3} \color{blue}{-3} }\) lets, just observe the \(\large\color{slate}{ f(x)=e^x }\) for now, though.

OpenStudy (solomonzelman):

1) When x gets infinitely small, (\(\large\color{slate}{ -10 }\) , \(\large\color{slate}{ -100 }\) , \(\large\color{slate}{ -1000 }\) .... \(\large\color{slate}{ -\infty }\)) THEN, what happens to the y? 2) When x gets infinitely large, (\(\large\color{slate}{ 10 }\) , \(\large\color{slate}{ 100 }\) , \(\large\color{slate}{ 1000 }\) .... \(\large\color{slate}{ +\infty }\)) THEN, what happens to the y?

OpenStudy (solomonzelman):

\(\large\color{blue}{ f(x)=e^{x} }\) Can you evaluate the \(\large\color{blue}{ f(x) }\) when x=-10 ?

OpenStudy (solomonzelman):

(you can just tell me the exact value)

OpenStudy (anonymous):

1/e^10?

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

So, for x=-100 and x=-1000, it would be, \(\large\color{blue}{ f(-100)=1/e^{100} }\) \(\large\color{blue}{ f(-1000)=1/e^{1000} }\)

OpenStudy (solomonzelman):

Which has a smaller value, \(\large\color{blue}{ 1/e^{100} }\) or \(\large\color{blue}{ 1/e^{1000} }\) ?

OpenStudy (anonymous):

1/e^100

OpenStudy (solomonzelman):

Which is smaller, 1/100 or 1/1000 ?

OpenStudy (anonymous):

1/1000

OpenStudy (anonymous):

that made a little more sense

OpenStudy (solomonzelman):

Wouldn't \(\large\color{blue}{ 1/{1000} }\) be smaller than \(\large\color{blue}{ 1/{100} }\) ?

OpenStudy (solomonzelman):

yes, and so, \(\large\color{blue}{ 1/e^{1000} }\) is smaller than \(\large\color{blue}{ 1/e^{100} }\)

OpenStudy (solomonzelman):

So this is what we have: \(\large\color{blue}{ f(x)=e^{x} }\) \(\large\color{blue}{ f(-10)=1/e^{10} }\) \(\large\color{blue}{ f(-100)=1/e^{100} }\) \(\large\color{blue}{ f(-1000)=1/e^{1000} }\) \(\large\color{blue}{ f(-10000)=1/e^{10000} }\)

OpenStudy (anonymous):

yeahh

OpenStudy (solomonzelman):

So, as x becomes a very big negative, like -100, -1000, -10000 and on, the denominator is increasing, and thus the f(x) is approaching zero closer and closer

OpenStudy (solomonzelman):

\(\large\color{blue}{ 1/e^{100} }\) is pretty much zero. But \(\large\color{blue}{ 1/e^{10000} }\) is much closer to zero. (it has even less value)

OpenStudy (solomonzelman):

So, as x approaches \(\large\color{blue}{ -\infty}\), the f(x) approaches \(\large\color{blue}{ 0 }\)

OpenStudy (solomonzelman):

is this making sense?

OpenStudy (solomonzelman):

(\(\large\color{blue}{ -\infty }\) means negative infinity)

OpenStudy (solomonzelman):

Now, as x becomes a very big number, where is the y value going?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\large\color{blue}{ f(x)=e^x }\) as x is, 10, 100, 1000 and on, is y increasing or decreasing?

OpenStudy (solomonzelman):

sorry, but I am asking if f(x) becomes very large or very small, as x values are growing?

OpenStudy (solomonzelman):

(what do you think ?)

OpenStudy (anonymous):

decreasing